# Kekulé Superconductivity in Twisted Magic Angle Bilayer Graphene

### v1

@@ III Order Parameter and Quantum Textures
Figure 5: Order-parameter amplitude $|\Delta|$ (left axis) and chemical potential $\mu$ (right axis) versus coupling strength $V$. The dashed line marks the flat-band bottom. For attractive $V\approx 40\,\mathrm{meV}$ (about three times the onset $V_{c}$), the system crosses into the BEC-like regime.
III.3 (3) Nematic Nature of Superconducting order
@@
Figure 6: (a) Figures show the density of states (DOS) versus energy $\omega$ (meV) for four different gap amplitudes, with curve labels $|\Delta|\simeq\,2.0,\,1.5,\,1.3$ meV. As the order parameter is reduced, the bottom of the characteristic U-shaped gap narrows towards a V-shape. Note that there is a finite zero bias conductance for the latter case. In (b) we plot the lowest positive Bogoliubov deGennes eigenvalue within the mini-Brillouin zone for two values of the order parameter. This demonstrates that for sufficiently small $\Delta$, a small Bogoliubov Fermi surface appears and the (V-shaped) DOS spectrum becomes gapless. (c). For larger $\Delta$ (still away from the BEC regime), the (U-shaped) spectrum is fully gapped.
III.4 (4) U to V-like transitions in tunneling
@@
In the trilayer case, where Kekulé superconductivity is similarly observed [18] the evolution of the conductance with doping is of particular interest. The tunneling spectrum which reflects the quasi-particle density of states [36], has a V-like shape (often with a “zero bias conductance”) when filling is in the regime closer to $\nu=-3$, while it transitions to a U-shaped spectrum, when $\nu$ is nearer 0.2-0.2. It is generally believed that these spectra, then, reflect a transition from a gapped superconductor to a nodal superconductor. Important to this analysis, one can associate the U regime with a shorter Landau Ginsberg coherence length which reflects [37] stronger attractive interactions. In this vein, while originally this U transition was speculated to relate to a transition all the way to a BEC regime, this has since been effectively refuted [38].
@@
Although these experiments address the tri-layer case, it is useful here to study the density of states (DOS) for our two-flat-band model, using the parameter set corresponding to Fig. 2. As shown in Fig. 6(a), for a superconducting order parameter of $|\Delta|\geq 1.5$ meV, the DOS exhibits a clear U-shape, indicating that the quasiparticle excitations are fully gapped.
@@
As the order parameter is reduced, the flat bottom of this U-shaped gap narrows, eventually transitioning into a smoothed V-shape with a finite DOS at zero frequency. We find this corresponds to the fact that there are touching points leading to a very small Bogoliubov Fermi surface contained within the quasi-particle dispersions for the “V case” and missing for the “U case”, as can be seen in Fig. 6(b) and Fig. 6(c)  [^5] Although their Bogoliubov Fermi surfaces are larger, there is some similarity here to work from Christos et al, Nat. Commun. 14, 7134 and Agterberg et al, Phys. Rev. Lett. 118, 127001. Also important, there is generally a finite conductance at zero energy (zero bias conductance) when the V shape is present. This constitutes a prediction from the present theory.
We can show that the V-like regime originates from the (very minimal) lack of inversion symmetry (${\bf q}\rightarrow-{\bf q}$) in the Bistritzer–MacDonald model for a single valley bandstructure. While the U shaped spectrum reflects the fact that the singlet form factor is nodeless, More precisely it is useful to consider the Bogoliubov dispersion in the single flat-band limit:
$E({\bf q})=\delta\epsilon({\bf q})\pm\sqrt{\bar{\epsilon}({\bf q})^{2}+\Delta^{2}\varphi({\bf q})^{2}}$
where $\delta\epsilon({\bf q})=[\epsilon({\bf q})-\epsilon(-{\bf q})]/2$ and $\bar{\epsilon}({\bf q})=[\epsilon({\bf q})+\epsilon(-{\bf q})]/2$ are derived from the dispersion of the partially filled flat band, $\epsilon({\bf q})$. Here $\delta\epsilon({\bf q})$ reflects the lack of inversion symmetry.
@@
From this expression, two regimes emerge. When $\Delta$ is large enough such that the condition $|\Delta\varphi({\bf q})|>|\delta\epsilon({\bf q})|$ holds for all $\bf q$}, the dispersion is necessarily gapped, leading to the U-shaped DOS. Conversely, when $\Delta$ is small enough that $|\Delta\varphi({\bf q})|\leq|\delta\epsilon({\bf q})|$ for some $\bf q$}, gapless points can develop. This results in a V-shaped DOS generally associated with a small Bogoliubov Fermi surface, reflecting the sharpness of the V. We anticipate the existence of a critical value, $\Delta_{c}$, that marks a transition from a fully gapped to a gapless superconducting state.
@@
Experimental Tests: Zero-bias Conductance. These observations lead to a very simple self consistency test of the theory. We consider a plot of the zero bias conductance as a function of temperature as compared with the temperature dependence of order parameter $\Delta$. This is shown in Fig. 7 and is inspired by a related plot (see their Fig 3e) for the moiré materials in Ref. [40]. One can speculate that here the V-shaped tunneling spectra may be qualitatively different from the classic $d$-wave behavior in the cuprates [41] where a plot $dI/dV\propto V$ vanishes exactly at zero bias. Throughout the cuprate family, however, there are some indications of a finite zero bias conductance but these are generally understood [42] as associated with a Dynes lifetime parameter in the tunneling density of states. What is different about the moiré systems, is that the finite zero-bias conductance appears order-parameter controlled, not merely lifetime-broadened. This suggests a superconducting state whose excitation spectrum remains partially gapless due to intrinsic features of the paired state itself, as illustrated in Fig. 7. Here this reflects the presence of an (albeit tiny) Bogoliubov Fermi surface.
@@
Figure 7: Density of states and order parameter versus temperature, computed at the interaction strength V=15V=15meV. As temperature increases, the order parameter decreases while the DOS increases. This behavior is consistent with the presence of a Bogoliubov Fermi surface and can be seen to be related to Figure 3e in Ref. [40] (see text), which addresses the zero-bias conductance.
IV Conclusions
@@
We have three central results concerning the twisted graphene family in this paper which we argue have some support from experiment: these are nematic order[34], U to V transitions[36] and large gap to $T_{c}$ ratios[9]. All these results are generic consequences of Kekulé superconducting order and symmetry properties of the single moiré valley (BM) model. At the heart of our paper is the assumption of intra-valley (Kekulé) pairing in which the major support comes from an observed correlation [18] in these moiré systems between the highest Kekulé peaks which are associated with the strongest pairing regime.
@@
We argue against the interpretation that Kekulé order acts as a pairing ”glue” since we know in superconductors which have a non-phononic, bosonic (e.g., magnetic) pairing mechanism, the superconductivity is strongest when the non-superconducting, alternative ordering is weakest[43]. Rather Kekulé order seems to be intertwined with the pairing machinery as distinct from the pairing mechanism. This order then leads to a PDW-like description of the superconductivity[19, 20, 44].
@@ A.1 Bistritzer–MacDonald Model and Symmetries
$H=v\Big(-i\nabla-\tfrac{\tau_{z}}{2}\mathbf{q}_{1}+\mathbf{M}\Big)\cdot\boldsymbol{\sigma}_{\theta/2}+\big[T(\mathbf{r})\tau^{+}+\text{h.c.}\big],$ (17)
where $v$ is the Fermi velocity of monolayer graphene, $\mathbf{M}$ denotes the $M$-point of the mini Brillouin zone, $\tau$ and $\sigma$ act on layer and sublattice indices, and $\boldsymbol{\sigma}_{\pm\theta/2}$ denote Pauli matrices rotated by angle $\pm\theta/2$. The tunneling matrix $T(\mathbf{r})$ is parameterized by two amplitudes $w_{AA}$ and $w_{AB}$,
@@
with $U(\mathbf{x})$ a unitary operator. Explicitly, the mirror symmetry and threefold rotation[45] are represented by $U_{M_{y}}=\sigma_{x}\tau_{x}$ and
$U_{C_{3}}(\mathbf{x})=\exp\!\left(\tfrac{2\pi i}{3}\,\hat{z}\cdot\boldsymbol{\sigma}\right)\exp\!\left[i\Big(2\mathbf{M}+\mathbf{b}_{1}\tfrac{\tau_{z}+1}{2}\Big)\cdot\mathbf{x}\right].$ (22)
@@ A.2 Flat band Approximation
Here, $H_{LL}$ is the Hamiltonian in the flat-band subspace, $H_{LH}$ couples the flat and remote bands, and $H_{HH}$ is the Hamiltonian within the remote-band subspace. After the transformation, the new Hamiltonian $\tilde{H}_{LL}$ acquires a correction of order $\Delta^{2}/E_{\text{band}}$. When this energy is much smaller than the bandwidth $W$ of a single flat band, the isolated-flat-band description of intra-valley superconductivity is well-justified. By contrast, in perfectly flat bands ($W\!\to\!0$) or for sufficiently large $\Delta$, remote-band effects must be included.
A.3 Quantum Textures
We show that the quantum texture contribution (arising from the BM bandstructure wavefunctions) is strictly smaller than 1, and in practice much smaller. It suffices to analyze the diagonal singlet form factor $\Lambda^{s}_{nn,L}(\mathbf{q})$, since the off–diagonal singlets and all triplet factors are smaller than $\Lambda^{s}_{nn,L}$.
(i) Sublattice selectivity from AB bonds. Because the superconducting order we consider is AB–bond–like, the form factor in Eq. 11 involves products of sublattice–polarized components after the $C_{2}T$ transformation; effectively, only half of the wavefunction components contribute. In the nonchiral limit the Bloch states are not fully sublattice polarized (their weight is distributed between A and B), and by normalization of the Bloch wavefunction this immediately suppresses $\Lambda_{nn,L}(\mathbf{q})$ below 1.
(ii) Lack of inversion in a single valley. In a single–valley model, the $\mathbf{q}\!\to\!-\mathbf{q}$ inversion symmetry is absent, so in general $u_{\mathbf{G},a;n}({\bf q})\neq u_{-\mathbf{G},a;n}(-{\bf q})$. Since Cooper pairing combines $(\mathbf{G},-\mathbf{G})$ harmonics, this asymmetry further reduces the form factor. Define the even/odd combinations
$f^{\pm}_{\mathbf{G},a;n}(\mathbf{q})=\frac{1}{\sqrt{2}}\bigl[u_{\mathbf{G},a;n}(\mathbf{q})\pm u_{-\mathbf{G},a;n}(-\mathbf{q})\bigr].$ (25)
With the inner product $\langle x|y\rangle\equiv\sum_{\mathbf{G}}x^{*}_{\mathbf{G}}y_{\mathbf{G}}$, the diagonal form factor can be written as
$\Lambda^{s}_{nn,L}(0,\mathbf{q})=\bigl\langle f^{+}_{a,n}(\mathbf{q})\,\big|\,f^{+}_{a,n}(\mathbf{q})\bigr\rangle-\bigl\langle f^{-}_{a,n}(\mathbf{q})\,\big|\,f^{-}_{a,n}(\mathbf{q})\bigr\rangle.$ (26)
If $u_{\mathbf{G}}\neq u_{-\mathbf{G}}$, then the “odd” overlap $\langle f^{-}_{a,n}(\mathbf{q})|f^{-}_{a,n}(-\mathbf{q})\rangle\neq 0$ and subtracts from the “even” overlap, reducing $|\Lambda_{nn,L}|$.
Consequently, $\Lambda^{s}_{nn,L}(\mathbf{q})<1$, and these two effects typically make it much smaller than 1.
A.4 A single flat-band Analysis
Solutions of the multiband gap equation show $\Delta/T_{c}\sim 10$ where the $\Delta$ is close to the ground state value (Fig. 4). To understand the origin of this large ratio, we consider superconducting gap for a a single flat band. This approximation is justified since the intraband matrix elements of the pairing form factor are much larger than the interband counterparts.
The Bogoliubov-de Gennes (BdG) Hamiltonian for a single flat-band model can be expressed in the Nambu basis as:
$H_{\text{single}}=\begin{pmatrix}\epsilon({\bf q})&\varphi({\bf q})\Delta\\ \varphi({\bf q})\Delta&-\epsilon(-{\bf q})\end{pmatrix}$ (27)
Here, $\epsilon(\mathbf{q})$ represents the dispersion of the partially-filled flat band, and $\varphi(\mathbf{q})$ is the pairing form factor projected onto this band (i.e., $\Lambda_{11}(\mathbf{q})$).
The Bogoliubov quasi-particle dispersion is then given by
$E_{\pm}({\bf q})=\delta\epsilon({\bf q})\pm E({\bf q}).$ (28)
where $\delta\epsilon=[\epsilon({\bf q})-\epsilon(-{\bf q})]/2$ and $E_{\mathbf{q}}=\sqrt{\epsilon(\mathbf{q})^{2}+\Delta^{2}\varphi(\mathbf{q})^{2}}$. In this expression, the $\delta\epsilon({\bf q})$ term directly reflects the breaking of inversion symmetry, while the radical term, $\sqrt{\bar{\epsilon}({\bf q})^{2}+\Delta^{2}\varphi({\bf q})^{2}}$, corresponds to the conventional Bogoliubov dispersion found in centrosymmetric superconductors.
In this Methods section, we primarily consider a sufficiently large enough $\Delta$ such that the quasi-particle dipersion is fully gapped. This corresponds to the U-shaped regime in the DOS in Fig. 8. In this limit, we can analyze the gap equation666the small asymmetric shift $\delta\epsilon$ appears in the argument of the hyperbolic tangent, which is negligible in the regimes considered and does not affect our conclusions.:
$\frac{1}{V}=\left\langle\frac{\varphi(\mathbf{q})^{2}}{2E_{\mathbf{q}}}\tanh\!\left(\frac{E_{\mathbf{q}}}{2T}\right)\right\rangle_{\text{mBZ}},$ (29)
where $E_{\mathbf{q}}=\sqrt{\epsilon(\mathbf{q})^{2}+\Delta^{2}\varphi(\mathbf{q})^{2}}$ is the quasiparticle energy and $\varphi(\mathbf{q})$ is the form factor projected onto the flat band (e.g., $\Lambda_{11}(\mathbf{q})$).
We consider the zero-temperature limit. In general, the superconducting order parameter $\Delta$ depends on the interaction strength $V=V_{0}({\bf q}=0)$. In the BM model (without strain effects) the bands are extremely flat so that the bandwidth $W\sim 0.5\,\text{meV}$. We observe that $\Delta>W$ is satisfied for most $V$ associated with a superconducting phase, where consistent with experiment, $\Delta\sim 1\,\text{meV}$.
In the limit $\Delta>W$ the zero-temperature gap can be approximated as
$\Delta\approx\frac{V}{2}\,\big\langle\varphi(\mathbf{q})\big\rangle_{\text{mBZ}}.$
Narrow-band limit. In the limit $2T_{c}>W$ linearizing Eq. (29) yields
$T_{c}\approx\frac{V}{4}\,\big\langle\varphi(\mathbf{q})^{2}\big\rangle_{\text{mBZ}}.$
Combining this with the zero-temperature result gives Eq. 13 in the main-text. This expression holds when the bandwidth $W$ is the smallest relevant energy scale. The ratio on the right-hand side of Eq. 13 is strictly larger than $2$. This can be seen as due to the absence of inversion symmetry in a single valley. For typical parameters, the right-hand side for the single flat-band form factor is $\approx 7.8$, if we let $\varphi({\bf q})=\Lambda_{11,L}({\bf Q}=\Gamma,{\bf q})$, close to the numerically observed ratio for $\Delta/T_{c}$.
Small $T_{c}$ limit. In the regime $T_{c}\ll W$ and at low filling ($\mu\ll W$), the factor $\tanh(\epsilon/2T_{c})$ is sensitive to $T_{c}$ only for $|\epsilon|\sim T_{c}$; otherwise $\tanh(\epsilon/2T_{c})\approx\pm 1$. Since $T_{c}/W\ll 1$, only a narrow energy window near the Fermi level is relevant for determining $T_{c}$. Accordingly, we approximate the dispersion near the Fermi level. Because we are at low filling, an expansion about the band bottom is appropriate: $\epsilon(\mathbf{q})\simeq q^{2}/2M-\mu,$ where $M$ is the effective mass obtained by expanding around the band minimum. In two dimensions this implies $d^{2}q=M\,d\epsilon\,d\theta.$ Numerically, $\varphi(\mathbf{q})$ saturates rapidly with $|\mathbf{q}|$, and we approximate it as angle-only, $\varphi(\mathbf{q})\to\varphi(\theta)$.
The linearized ($\Delta\to 0$) gap equation becomes
$\frac{1}{V}\;\simeq\;\big\langle\varphi(\theta)^{2}\big\rangle_{\text{mBZ}}\,\frac{1}{W}\int_{0}^{W}\!d\epsilon\;\frac{1}{2\epsilon}\,\tanh\!\Big(\frac{\epsilon}{2T_{c}}\Big).$ (30)
Using the standard asymptotic
$\int_{0}^{W}\frac{d\epsilon}{\epsilon}\,\tanh\!\Big(\frac{\epsilon}{2T_{c}}\Big)=\ln\!\Big(\frac{c\,W}{T_{c}}\Big),$ (31)
with $c\approx 1.13$, Eq. (30) gives
$T_{c}\;\simeq\;c\,W\exp\!\Big[-\,\frac{2W}{V\langle\varphi^{2}\rangle_{\text{mBZ}}}\Big].$ (32)
Using the zero temperature result of order parameter, one finds
$\frac{\Delta}{T_{c}}\simeq\frac{V}{2cW}\,\big\langle\varphi\big\rangle_{\text{mBZ}}\,\exp\!\Big[\frac{2W}{V\langle\varphi^{2}\rangle_{\text{mBZ}}}\Big].$ (33)
We use typical parameters: $V=10~\mathrm{meV}$,  $\langle\varphi\rangle_{\text{mBZ}}=0.24$,  $\langle\varphi^{2}\rangle_{\text{mBZ}}=0.063$,  $W=0.5~\mathrm{meV}$,  and Eq. 33 is evaluated to be $\Delta/T_{c}\simeq 10$.
@@ A.4 A single flat-band Analysis
Figure 8: U-shaped density of states (DOS) from two-flat-band superconductivity. The figure is plotted at $|\Delta|\simeq 2.5~\mathrm{meV}$. From the DOS, the coherence peak appears at $\omega=\Delta_{t}\simeq 1~\mathrm{meV}$, which is smaller than the order parameter $|\Delta|$. The dip feature around $\omega\sim$ 1meV originates from the band touching of two BM flat bands at $K/K^{\prime}$ points.

### v2

@@ III.1 Establishing the Thermodynamically Stable Phase
Figure 4: Order-parameter amplitude $|\Delta|$ (left axis) and chemical potential $\mu$ (right axis) versus coupling strength $V$. The dashed line marks the flat-band bottom. For attractive $V\approx-1.7\,\mathrm{meV}$ ($\Delta\simeq 0.73$meV), the system crosses into the BEC-like regime.
III.2 Nematic Nature of Superconducting order
@@
Figure 5: (a) Density of states (DOS) versus energy $\omega$ (meV) for three different gap amplitudes, with curves labeled by $|\Delta|\simeq 0.49,0.36,$ and $0.32$ meV. As the order parameter is reduced, the bottom of the characteristic U-shaped gap narrows towards a V-shape. Notably, a finite zero-bias conductance emerges in the latter case (b) The lowest positive Bogoliubov-de Gennes eigenvalue $E_{1}(M,\mathbf{k})$ plotted across the mini-Brillouin zone for $|\Delta|=0.49$ meV. At this amplitude (still within the BCS-like regime), the spectrum remains fully gapped, corresponding to the U-shaped DOS. (c). At a smaller amplitude of $|\Delta|=0.32$ meV, a small Bogoliubov Fermi surface appears and the (V-shaped) DOS spectrum becomes gapless.
IV U to V-like transitions in tunneling
@@
In the trilayer case, where Kekulé superconductivity is similarly observed [Kim2023] the evolution of the conductance with doping is of particular interest. The tunneling spectrum which reflects the quasi-particle density of states [Kim2022], has a V-like shape (often with a “zero bias conductance”) when filling is in the regime closer to $\nu=-3$, while it transitions to a U-shaped spectrum, when $\nu$ is nearer 0.2-0.2. It is generally believed that these spectra, then, reflect a transition from a gapped superconductor to a nodal superconductor. Important to this analysis, one can associate the U regime with a shorter Landau Ginsberg coherence length which reflects [Chen2024] stronger attractive interactions. In this context, while originally this U transition was speculated to relate to a transition all the way to a BEC regime, this may have been refuted [Van2025].
@@
Although these experiments address the tri-layer case, it is useful here to study the density of states (DOS), using the parameter set corresponding to Fig. 2. As shown in Fig. 5(a), for a superconducting order parameter of $|\Delta|\geq 0.4$ meV, the DOS exhibits a clear U-shape, indicating that the quasiparticle excitations are fully gapped.
@@
As the order parameter is reduced, the flat bottom of this U-shaped gap narrows, eventually transitioning into a smoothed V-shape with a finite DOS at zero frequency. We find this corresponds to the fact that there are touching points leading to a very small Bogoliubov Fermi surface contained within the quasi-particle dispersions for the “V case” and missing for the “U case”, as can be seen in Fig. 5(b) and Fig. 5(c)  [^8] Although their Bogoliubov Fermi surfaces are larger, there is some similarity here to work from Christos et al, Nat. Commun. 14, 7134 and Agterberg et al, Phys. Rev. Lett. 118, 127001. We stress here that there is generally a finite conductance at zero energy (zero bias conductance) when the V shape is present. This constitutes a prediction from the present theory.
To address this in more detail, it is useful to consider the Bogoliubov dispersion in the two-flat-band limit:
$E(M,{\bf q})=\delta\epsilon({M,\bf q})\pm\sqrt{\bar{\epsilon}({M,\bf q})^{2}+\Delta^{2}\varphi({M,\bf q})^{2}}$ (10)
where $\delta\epsilon({M,\bf q})=[\epsilon_{1}(M+{\bf q})-\epsilon_{2}(M-{\bf q})]/2$ and $\bar{\epsilon}({M,\bf q})=[\epsilon_{1}(M+{\bf q})+\epsilon_{2}(M-{\bf q})]/2$ are derived from the dispersion of two flat bands, $\epsilon_{1,2}({M,\bf q})$. Here $\varphi(M,q)=\sum_{L}\Lambda^{t}_{12,L}(M,q)$ represents the strength of off-diagonal pairing [^9] From Fig. 2, we observe that the off-diagonal components of $\Lambda$ are significantly larger than the diagonal ones. Consequently, we retain only the off-diagonal pairing..
@@
From this expression, two regimes emerge. When $\Delta$ is large enough such that the condition $|\Delta\varphi(M,{\bf q})|>|\delta\epsilon(M,{\bf q})|$ holds for all $\bf q$}, the dispersion is necessarily gapped, leading to the U-shaped DOS. Conversely, when $\Delta$ is small enough that $|\Delta\varphi(M,{\bf q})|\leq|\delta\epsilon(M,{\bf q})|$ for some $\bf q$}, gapless points can develop. This results in a V-shaped DOS generally associated with a small Bogoliubov Fermi surface, reflecting the sharpness of the V. We anticipate the existence of a critical value, $\Delta_{c}$, that marks a transition from a fully gapped to a gapless superconducting state.
V Zero-bias Conductance
Of interest is the behavior of the zero-bias conductance in these tunneling spectra. The fact that two different groups [Park2025, Kim2022] have arrived at quite similar plots suggests that this may be an intrinsic phenomenon. What is important here is that the zero bias conductance, in principle, contains information about the excitation spectrum of the superconductor. Moreover, there should be some distinction between the U and V shaped cases which will be important to assess in future, although the experiments thus far have been in the V-shaped regime.
We present a plot of the zero bias conductance in this regime, as a function of temperature which is compared with the temperature dependence of order parameter $\Delta$. This is shown in Fig. 6 and is inspired by a related plot (see their Fig 3e) for the moiré materials in Ref. [Park2025]. In the Methods section we discuss the U-shaped counterpart.
@@ V Zero-bias Conductance
One can speculate that here the V-shaped tunneling spectra may be qualitatively different from the classic $d$-wave behavior in the cuprates [Pushp2009] where a plot $dI/dV\propto V$ vanishes exactly at zero bias. Throughout the cuprate family, however, there are some indications of a finite zero bias conductance but these are generally understood [Reber2013] as associated with a Dynes lifetime parameter in the tunneling density of states. We argue that what is different about the moiré systems is that the finite zero-bias conductance appears order-parameter controlled, not merely lifetime-broadened. This suggests a superconducting state whose excitation spectrum remains partially gapless due to intrinsic features of the paired state itself, as illustrated in Fig. 6. Here the theory for the graphene case reflects the presence of a Bogoliubov Fermi surface. Given the consistency across a few fillings  [^10] Private Communications with Hyunjin Kim and across groups [Park2025, Kim2022], we attribute the finite zero-bias conductance to an intrinsic Bogoliubov Fermi surface rather than to extrinsic, random Dynes-like broadening.
@@
Figure 6: Density of states and order parameter versus temperature, computed at the interaction strength V=0.8V=-0.8meV. As temperature increases, the order parameter decreases while the DOS increases. This behavior is consistent with the presence of a Bogoliubov Fermi surface and can be seen to be related to Figure 3e in Ref. [Park2025] (see text), which addresses the zero-bias conductance.
VI Conclusions
@@
In addition to these zero bias predictions, we have three central results concerning the twisted graphene family in this paper which we argue have some support from experiment: these are nematic order [caoNematic], U to V transitions [Kim2022] and triplet pairing [Cao2021]. All these results are generic consequences of Kekulé superconducting order and symmetry properties of the single moiré valley (BM) model. At the heart of our paper is the assumption of intra-valley (Kekulé) pairing in which the major support comes from an observed correlation [Kim2023] in these moiré systems between the highest Kekulé peaks which are associated with the strongest pairing regime.
@@
We argue against the interpretation that Kekulé order acts as a pairing ”glue” since we know in superconductors which have a non-phononic, bosonic (e.g., magnetic) pairing mechanism, the superconductivity is strongest when the non-superconducting, alternative ordering is weakest[zq2025]. Rather Kekulé order seems to be intertwined with the pairing machinery as distinct from the pairing mechanism. This order then leads to a PDW-like description of the superconductivity[Tsuchiya2016, Roy2010, PhysRevLett.114.237001, wang2025].
@@ A.1 Bistritzer–MacDonald Model and Symmetries
$H=v\Big(-i\nabla-L_{z}\mathbf{q}_{1}/2+\mathbf{M}\Big)\cdot\boldsymbol{\sigma}_{\theta/2}+\big[T(\mathbf{r})L^{+}+\text{h.c.}\big],$ (13)
where $v$ is the Fermi velocity of monolayer graphene, $\mathbf{M}$ denotes the $M$-point of the mini Brillouin zone, $L$ and $\sigma$ act on layer and sublattice indices, and $\boldsymbol{\sigma}_{\pm\theta/2}$ denote Pauli matrices rotated by angle $\pm\theta/2$. The tunneling matrix $T(\mathbf{r})$ is parameterized by two amplitudes $w_{AA}$ and $w_{AB}$,
@@
with $U(\mathbf{x})$ a unitary operator. Explicitly, the mirror symmetry and threefold rotation[PhysRevB.99.035111] are represented by $U_{M_{y}}=\sigma_{x}L_{x}$ and
$U_{C_{3}}(\mathbf{x})=\exp\!\left(\tfrac{2\pi i}{3}\,\hat{z}\cdot\boldsymbol{\sigma}\right)\exp\!\left[i\Big(2\mathbf{M}+(L_{z}+1)\mathbf{b}_{1}/2\Big)\cdot\mathbf{x}\right].$ (18)
@@ A.2 Flat band Approximation
Here, $H_{LL}$ is the Hamiltonian in the flat-band subspace, $H_{LH}$ couples the flat and remote bands, and $H_{HH}$ is the Hamiltonian within the remote-band subspace. After the transformation, the new Hamiltonian $\tilde{H}_{LL}$ acquires a correction of order $\Delta^{2}/E_{\text{band}}$. When this energy is much smaller than the bandwidth $W$ of a single flat band, the isolated-flat-band description of intra-valley superconductivity is well-justified, provided the bands are not perfectly flat ($W\!\to\!0$) or $\Delta$ sufficiently large so that remote-band effects must be included.
In numerical simulations that incorporate remote bands into the gap equation, we find that including only the two flat bands yields reasonable results, although adding more bands enhances precision. In practice, we retain $10$ bands.
@@
Figure 7: U-shaped density of states (DOS) from two-flat-band superconductivity. The figure is plotted at $|\Delta|\simeq 0.63~\mathrm{meV}$ ($V\simeq-1.5$ meV). The dip feature around $\omega\sim$ 1meV originates from the band touching of two BM flat bands at $K/K^{\prime}$ points.
Figure 8: Quasi-particle density of states (reflecting zero bias conductance) and order parameter versus temperature, computed at the interaction strength $V\simeq-1.17$meV. This is in the U-shaped tunneling regime and should be contrasted with Fig. 6 in the main text which is for the V-shaped case with weaker attractive interaction. As temperature increases, the density of states remains gapped until the order parameter $\Delta$ drops sufficiently and then the system crosses over to a more V-like spectrum associated with a Bogoliubov Fermi surface.